# From minimal-length quantum theory to modified gravity

### v1

@@ III.1 Extension to $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity
where we define $P\equiv R_{\mu\nu}R^{\mu\nu}$. The inclusion of such terms is strongly motivated by various independent lines of reasoning. From a quantum field theoretic perspective, one-loop effective actions of matter fields propagating in curved backgrounds inevitably generate curvature-squared contributions, including both $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$, with coefficients fixed by the matter content and regularization scheme employed [Birrell:1982ix, Parker:2009uva]. These terms also arise naturally in the context of renormalization group flows of gravitational couplings, where they play a central role in the construction of asymptotically safe theories of gravity [Reuter:1996cp].
Furthermore, from the viewpoint of higher-derivative gravity, the addition of the $R_{\mu\nu}R^{\mu\nu}$ invariant alters the spectrum of gravitational excitations, introducing additional degrees of freedom that can improve the ultraviolet behavior of the theory and potentially render it perturbatively renormalizable [Stelle:1976gc]. Together, these arguments provide compelling evidence that effective gravitational theories incorporating quantum corrections should generically depend on both $R$ and $P$, with implications for both classical dynamics and BH thermodynamics.
Wald’s entropy formula for a generic action depending on arbitrary curvature invariants reads [DAgostino:2024sgm, DAgostino:2024ymo]
@@ 2 $\ln A$ $R^{2}\left(\ln R^{2}-1\right)$
Notably, the appearance of both $\sqrt{A}$ and $\ln A$ corrections is in agreement with a wide range of results found in the literature [Carlip:2000nv, Kaul:2000kf, Sen:2012dw, Adler:2001vs, Medved:2004yu, Nozari:2008rc, Majumder:2012rtc, Tawfik:2015kga].
@@
The presence of the $R^{3/2}$ and $R^{2}\ln R^{2}$ terms is of particular interest. The former reflects the impact of linear GUP corrections, while the latter combines the effects of both linear and quadratic deformations. Such non-analytic curvature contributions often emerge in quantum corrections to the effective action, e.g., from trace anomalies or conformal field theories on curved backgrounds [Duff:1993wm, Birrell:1982ix, Avramidi:2000bm], as well as from gravity models with entropy-based reconstruction schemes [Cognola:2005de]. This further supports the idea that GUP-inspired corrections may capture essential features of an underlying quantum theory of gravity.
@@ V Astrophysical application
As an astrophysical application of the framework developed above, we now consider an extended gravity theory inspired by the Lagrangian (50), where we restore the radial dependence of the Ricci scalar $R$. For this purpose, by varying the action with respect to the metric tensor, we obtain the $f(R)$ field equations [Sotiriou:2008rp]:
@@
In the last step of Eq. (54), we have neglected the $R^{2}\ln R^{2}$ term, which is sub-leading with respect to the $R^{2}$ term at astrophysical scales [^3] We remind that Eq. (54) is up to an integration constant that appears as $\epsilon R^{2}\ln\mu$, where $\mu$ has dimensions of $R^{-2}$. In our context, $\mu$ can be fixed such that $\mu R^{2}\sim 1$ at astrophysical scales.. This framework corresponds to the well-known Starobinsky model [Starobinsky:1980te], which reproduces GR in the limit where the GUP reduces to the Heisenberg principle, namely for $c_{2}\rightarrow 0$.
We can attempt to solve the resulting differential problem in the weak-field slow-motion limit of the theory [Schmidt:1986via, Capozziello:2009zz, Berry:2011pb]. To do that, we search for solutions under the following metric ansatz:
@@
The above equation represents a fourth-order gravity theory motivated by quantum considerations. Therefore, the solutions must guarantee a Newtonian behavior in the long-range regime, i.e., they should be exponentially vanishing [Stelle:1977ry]. In particular, taking $\rho$ as a delta source in Eq. (61), i.e. $\rho=M\delta$, we get the Yukawa-like solutions
@@
with $\epsilon>0$ to avoid oscillating behaviors. In light of Eq. (55), this condition implies $c_{2}>0$. From Eq. (53), this is precisely the requirement for the existence of a minimal length [Kempf:1994su]. It is straightforward to verify that Eqs. (62)–(63) reproduce the linearized Schwarzschild metric in the regime $r\gg\sqrt{12\epsilon}$, i.e., at distances well above the Planck scale.
We can now examine the observable consequences of the GUP-corrected metric by analyzing the light deflection by a massive body. To this aim, we follow the treatment of [Weinberg:1972kfs, Scardigli:2014qka] and consider a photon approaching the Sun from very large distances. Then, the photon orbit $\vartheta(r)$ in given by
@@
which coincides with the prediction of the Schwarzschild metric arising from the second-order expansion in $GM/r_{0}\ll 1$  [^5] In the literature, Eq. (75) is often expressed as [Epstein:1980dw, Fischbach:1980su, Richter:1982zz] $\Delta\vartheta\approx\frac{4GM}{b}+\frac{15\pi}{4}\left(\frac{GM}{b}\right)^{2},$ (76) where $b=r_{0}/\sqrt{1-2GM/r_{0}}$ is the impact parameter. One can readily check that Eqs. (75) and (76) agree up to second order in $GM/r_{0}\ll 1$..
@@
where $GM_{\odot}\approx 7.5\times 10^{9}\ \text{eV}^{-1}$ and $R_{\odot}\approx 3.5\times 10^{15}\,\text{eV}^{-1}$. Here, $\gamma$ is the standard post-Newtonian parameter, whose value has been measured, e.g., with the very long baseline interferometry (VLBI) [Lambert:2009xy]:
@@
Finally, a rough estimate of the quadratic GUP correction (53) follows from Eq. (55), from which we get $c_{2}=16\epsilon\zeta^{2}G^{-1}$. Recalling that $\zeta=\Delta x\,A^{-1/2}$ (see below Eq. (8)) and assuming $\Delta x\sim\ell_{P}$, we find $c_{2}\sim 16\epsilon A^{-1}$. Since in our framework $A$ must be interpreted as the horizon area of a solar-mass BH, i.e., $A=4\pi R_{S,\odot}^{2}$ with $R_{S,\odot}\approx 1.5\times 10^{10}\,\text{eV}^{-1}$, we finally obtain
@@
From Eq. (53), we know that $\Delta x_{\text{min}}=\sqrt{c_{2}}\ell_{P}$ [Kempf:1994su], we get $\Delta x_{\text{min}}\lesssim 9.2\times 10^{3}\ell_{P}$. It is worth emphasizing that the constraint (80) provides a stringent limit on the quadratic GUP model, improving upon previous results in the literature (see, e.g., [Bosso:2023aht] and references therein).
@@ Acknowledgments
R.D. acknowledges financial support from INFN-Sezione di Roma 1, esperimento Euclid. The research of G.G.L. is supported by the Ayuda Postdoctoral of the University of Lleida (project code: X25027). G.G.L. gratefully acknowledges the contribution from Cost Action CA21106 – COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers). P.B. and G.G.L. acknowledge the contribution of the LISA Cosmology Working Group (CosWG), as well as support from Cost Action CA23130 – Bridging high and low energies in search of quantum gravity (BridgeQG). R.D. and G.G.L. acknowledge work from COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by European Cooperation in Science and Technology.

### v2

@@ III.1 Extension to $f(R,R_{\mu\nu}R^{\mu\nu})$ gravity
where we define $P\equiv R_{\mu\nu}R^{\mu\nu}$. The inclusion of such terms is strongly motivated by various independent lines of reasoning. From a quantum field theoretic perspective, one-loop effective actions of matter fields propagating in curved backgrounds inevitably generate curvature-squared contributions, including both $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$, with coefficients fixed by the matter content and regularization scheme employed Birrell and Davies (1982); Parker and Toms (2009). These terms also arise naturally in the context of renormalization group flows of gravitational couplings, where they play a central role in the construction of asymptotically safe theories of gravity Reuter (1998).
Furthermore, from the viewpoint of higher-derivative gravity, the addition of the $R_{\mu\nu}R^{\mu\nu}$ invariant alters the spectrum of gravitational excitations, introducing additional degrees of freedom that can improve the ultraviolet behavior of the theory and potentially render it perturbatively renormalizable Stelle (1977). Together, these arguments provide compelling evidence that effective gravitational theories incorporating quantum corrections should generically depend on both $R$ and $P$, with implications for both classical dynamics and BH thermodynamics.
Wald’s entropy formula for a generic action depending on arbitrary curvature invariants reads D’Agostino and Luciano (2024); D’Agostino et al. (2024)
@@ 2 $\ln A$ $R^{2}\left(\ln R^{2}-1\right)$
Notably, the appearance of both $\sqrt{A}$ and $\ln A$ corrections is in agreement with a wide range of results found in the literature Carlip (2000); Kaul and Majumdar (2000); Sen (2013); Adler et al. (2001); Medved (2005); Nozari and Mehdipour (2008); Majumder (2013); Tawfik and El Dahab (2015).
@@
The presence of the $R^{3/2}$ and $R^{2}\ln R^{2}$ terms is of particular interest. The former reflects the impact of linear GUP corrections, while the latter combines the effects of both linear and quadratic deformations. Such non-analytic curvature contributions often emerge in quantum corrections to the effective action, e.g., from trace anomalies or conformal field theories on curved backgrounds Duff (1994); Birrell and Davies (1982); Avramidi (2000), as well as from gravity models with entropy-based reconstruction schemes Cognola et al. (2005). This further supports the idea that GUP-inspired corrections may capture essential features of an underlying quantum theory of gravity.
@@ V Astrophysical application
As an astrophysical application of the framework developed above, we now consider an extended gravity theory inspired by the Lagrangian (50), where we restore the radial dependence of the Ricci scalar $R$. For this purpose, by varying the action with respect to the metric tensor, we obtain the $f(R)$ field equations Sotiriou and Faraoni (2010):
@@
In the last step of Eq. (54), we have neglected the $R^{2}\ln R^{2}$ term, which is sub-leading with respect to the $R^{2}$ term at astrophysical scales [^4] We remind that Eq. (54) is up to an integration constant that appears as $\epsilon R^{2}\ln\mu$, where $\mu$ has dimensions of $R^{-2}$. In our context, $\mu$ can be fixed such that $\mu R^{2}\sim 1$ at astrophysical scales.. This framework corresponds to the well-known Starobinsky model Starobinsky (1980), which reproduces GR in the limit where the GUP reduces to the Heisenberg principle, namely for $c_{2}\rightarrow 0$.
We can attempt to solve the resulting differential problem in the weak-field slow-motion limit of the theory Schmidt (1986); Capozziello et al. (2009); Berry and Gair (2011). To do that, we search for solutions under the following metric ansatz:
@@
The above equation represents a fourth-order gravity theory motivated by quantum considerations. Therefore, the solutions must guarantee a Newtonian behavior in the long-range regime, i.e., they should be exponentially vanishing Stelle (1978). In particular, taking $\rho$ as a delta source in Eq. (61), i.e. $\rho=M\delta$, we get the Yukawa-like solutions
@@
with $\epsilon>0$ to avoid oscillating behaviors. In light of Eq. (55), this condition implies $c_{2}>0$. From Eq. (53), this is precisely the requirement for the existence of a minimal length Kempf et al. (1995). It is straightforward to verify that Eqs. (62)–(63) reproduce the linearized Schwarzschild metric in the regime $r\gg\sqrt{12\epsilon}$, i.e., at distances well above the Planck scale.
We can now examine the observable consequences of the GUP-corrected metric by analyzing the light deflection by a massive body. To this aim, we follow the treatment of Weinberg (1972); Scardigli and Casadio (2015) and consider a photon approaching the Sun from very large distances. Then, the photon orbit $\vartheta(r)$ in given by
@@
which coincides with the prediction of the Schwarzschild metric arising from the second-order expansion in $GM/r_{0}\ll 1$  [^6] In the literature, Eq. (75) is often expressed as Epstein and Shapiro (1980); Fischbach and Freeman (1980); Richter and Matzner (1982) $\Delta\vartheta\approx\frac{4GM}{b}+\frac{15\pi}{4}\left(\frac{GM}{b}\right)^{2},$ (76) where $b=r_{0}/\sqrt{1-2GM/r_{0}}$ is the impact parameter. One can readily check that Eqs. (75) and (76) agree up to second order in $GM/r_{0}\ll 1$..
@@
where $GM_{\odot}\approx 7.5\times 10^{9}\ \text{eV}^{-1}$ and $R_{\odot}\approx 3.5\times 10^{15}\,\text{eV}^{-1}$. Here, $\gamma$ is the standard post-Newtonian parameter, whose value has been measured, e.g., with the very long baseline interferometry (VLBI) Lambert and Le Poncin-Lafitte (2009):
@@
Finally, a rough estimate of the quadratic GUP correction (53) follows from Eq. (55), from which we get $c_{2}=16\epsilon\zeta^{2}G^{-1}$. Recalling that $\zeta=\Delta x\,A^{-1/2}$ (see below Eq. (8)) and assuming $\Delta x\sim\ell_{P}$, we find $c_{2}\sim 16\epsilon A^{-1}$. Since in our framework $A$ must be interpreted as the horizon area of a solar-mass BH, i.e., $A=4\pi R_{S,\odot}^{2}$ with $R_{S,\odot}\approx 1.5\times 10^{10}\,\text{eV}^{-1}$, we obtain
@@
Since $\Delta x_{\text{min}}=\sqrt{c_{2}}\ell_{P}$ Kempf et al. (1995), from Eq. (53) we finally get
$\Delta x_{\text{min}}\lesssim 9.2\times 10^{3}\ell_{P}\,$, (81)
which significantly improves upon previous upper bounds in the literature (see, e.g., Ref. Bosso et al. (2023a) and references therein).
V.1 Comparison with previous results
It is now instructive to compare our analysis and results with related studies reported in the literature. For instance, Ref. Hammad (2015) investigates the connection between $f(R)$-modified gravity and generalized uncertainty relations via Wald’s entropy, but restricts the analysis to the quadratic GUP model.
A different strategy is adopted in Ref. Ali and Wojnar (2024), where a covariant framework for a linear GUP, together with a corresponding deformation of the Poincaré algebra, is formulated within the context of metric-affine gravity. While this approach establishes links with Ricci-based models and explores implications for the covariant entropy bound, it does not provide a systematic method for reconstructing higher-curvature gravitational actions from arbitrary GUP deformations.
Similarly, Ref. Dias et al. (2016) relates the GUP to extended theories of gravity only at the level of linearized weak-field dynamics. Such a procedure, however, does not lead to a covariant gravitational action nor guarantee compatibility with the thermodynamic identities underlying BH entropy.
By contrast, in the present work, we start from a general form of the GUP, derive the corresponding entropy correction in closed form, and then employ Wald’s Noether-charge formula to reconstruct a fully covariant extended-gravity action. The theoretical framework developed here is therefore substantially more general and robust, as it establishes a direct and systematic correspondence between minimal-length quantum mechanics and gravitational dynamics at the level of the full action.
It is also interesting to compare our numerical constraint with the experimental bounds obtained within the approach of Ref. Dias et al. (2016). In particular, the torsion-based constraint reported there yields $\sqrt{\beta_{g}}<2.3\times 10^{-5}\,\text{m}$, which corresponds to $\Delta x_{\text{min}}\lesssim 1.4\times 10^{30}\,\ell_{P}$. Comparing this result with the upper bound (81), we find that our constraint is roughly 26 orders of magnitude tighter. Moreover, even considering the most stringent bound reported in Ref. Dias et al. (2016), inferred from the electron anomalous magnetic moment, namely $\Delta x_{\text{min}}\lesssim 4.7\times 10^{-18}\,\text{m}=2.9\times 10^{17}\,\ell_{P}$, our result still represents an improvement by approximately 13 orders of magnitude.
To summarize, we point out that the upper bound (81) obtained in this work is not only more stringent than the aforementioned ones, but it is among the most stringent results reported so far, as one can easily check by comparing the bounds collected in Ref. Bosso et al. (2023a).
@@ Acknowledgments
R.D. acknowledges financial support from Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Roma 1, esperimento Euclid. The research of G.G.L. is supported by the Ayuda Postdoctoral of the University of Lleida (project code: X25027). G.G.L. gratefully acknowledges the contribution from Cost Action CA21106 – COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers). P.B. and G.G.L. acknowledge the contribution of the LISA Cosmology Working Group (CosWG), as well as support from Cost Action CA23130 – Bridging high and low energies in search of quantum gravity (BridgeQG). R.D. and G.G.L. acknowledge work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by European Cooperation in Science and Technology.